what is 100^2 and 1000^2
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

100^2 = 100 * 100 = 1 * 10^2 * 1 * 10^2 = 1 * 10^4 = 10,000

1000^2 = 1000 * 1000 = 1 * 10^3 * 1 * 10^3 = 1 * 10^6 = 1,000,000.

with exponents, the following rules apply.

x^2 * x^2 = x^(2+2) = x^4

x^3 * x^3 = x^(3+3) = x^6

here's a tutorial on exponent arithmetic

https://www.rapidtables.com/math/number/exponent.html

Answer by MathTherapy(10837)

what is 100^2 and 1000^2
What are you asking?
Normally "and" means SUM or ADD! Did you mean: 1002 + 1,0002, or 1002 * 1,0002?